can you help me solve 10^6x = 93 and round to the nearest ten-thousandths please? I don't have a graphic calculator and don't understand it.
Answer by htmentor(1343)

Take the log10 of both sides:
log(10^6x) = log(93)
6x*log10 = log(93)
x = log(93)/6 = 0.3281